OpenStudy (sushi121212):

Factor: 8c^2+25ck+3k^2

OpenStudy (3mar):

If I may help?

OpenStudy (sushi121212):

Yes, please!!!

OpenStudy (bonnieisflash1.0):

need help?

OpenStudy (sushi121212):

Yes, please!! I don't understand this.

OpenStudy (3mar):

\[8c^2+25ck+3k^2=\huge (8c+1)(c+3)\] Does it make sense?

OpenStudy (sushi121212):

What about the k?

OpenStudy (bonnieisflash1.0):

@3mar you can take this question

OpenStudy (3mar):

What do you think? @Sushi121212 \[8c^2+25ck+3k^2=\huge (8c+1)(c+3)\] \[8c^2+25ck+3k^2=\huge (8c+k)(c+3k)\] so what is the value of k?

OpenStudy (3mar):

Conclusion: \[\huge k=1\]